# New Linear Model of a Composite Energy Storage System with Realizable Dispatch Guarantees

### v1

@@ III-D Realizable Composite Power Trajectories
If $\Delta E[0]\leq\Delta E_{\text{max}}$, then any feasible composite battery dispatch schedule $\{P_{\text{c/d}}^{\ast}[k]\}_{k=0}^{K-1}$ resulting from optimizing over (12) will satisfy (1) under PSC and is realizable across $N$ elements.
@@
$\frac{P_{\text{c}}[k]}{N{P}_{\text{c,max}}}$ $+\frac{P_{\text{d}}[k]}{N{P}_{\text{d,max}}}\leq\frac{N-1}{N}$ (12b)
$N\Delta E_{\text{max}}$ $\leq E[k]\leq N\left(E_{\text{max}}-\Delta E_{\text{max}}\right)$ (12c)
@@
From Proposition 1, (12b) ensures satisfaction of (1c). Similarly, with $\Delta E[0]\leq\Delta E_{\text{max}}$, Proposition 2 guarantees that (1d) holds for all $N$ elements. Thus, any feasible composite dispatch is realizable by its $N$ elements.∎
Note that for a fixed composite battery, if the elements’ $P_{\text{c/d,max}}$ decreases (i.e., $N$ increases), the cutting plane in (12b) approaches the cutting plane of the relaxed model. Furthermore, for $M\gg 1$, $\Delta E_{\text{max}}$ tends to zero. That is, with faster control and more granular elements, (12) approaches the relaxed battery model defined by (12a); however, with guaranteed realizability of any feasible dispatch.
@@ IV Simulation Results
We apply the realizable composite battery model (RCB) in (12) to minimize power tracking error for a given reference signal based on the objective function $\sum_{k=0}^{K}\left((P_{\text{c}}[k]-P_{\text{d}}[k])-P_{\text{ref}[k]}\right)^{2}$. For comparison, relaxed and robust models from [RobustNawaf] are used with equal power-sharing enforced. Also, a Mixed-Integer Linear Program (MILP) formulation with unequal power-sharing for $N$ elements in (1) is used for comparison. The chosen element battery parameters are obtained from a Tesla Powerwall [tesla_2019] and we set $N=1000$ and $M=1$. The results from using each model are shown in Fig. 3.
@@
The RCB model and relaxed battery model produce the same optimal dispatch. However, this optimal dispatch was not realizable under equal power sharing (relaxed) causing the battery to saturate and cause the sudden drop in power seen in Fig. 3. The MILP, robust, and RCB models produced realizable results, with the robust formulation resulting in worse tracking performance. The MILP and RCB formulations produced comparable results, however, with a significant computational advantage for the RCB model.
@@ V Conclusion
A new convex dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements to realize simultaneous charging and discharging. A priority-based disaggregation strategy is described and used to characterize sufficient conditions on the composite battery dispatch under which element constraints are guaranteed to be satisfied. Previous battery optimization implicitly assumed equal power sharing between all elements. By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models, which yield realizable power trajectories.

### v2

@@ III-D Realizable Composite Power Trajectories
If $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$, then any feasible composite battery dispatch schedule $\{P_{\text{c}}[k],P_{\text{d}}[k]\}_{k=0}^{K-1}$ satisfying
@@
$\frac{P_{\text{c}}[k]}{N{P}_{\text{c,max}}}+\frac{P_{\text{d}}[k]}{N{P}_{\text{d,max}}}$ $\leq\frac{N-1}{N}$ $\forall k=0,\ldots,K-1$ (12b)
$N\varepsilon\leq E[k]$ $\leq N\left(E_{\text{max}}-\varepsilon\right)$ $\forall k=0,\ldots,K$ (12c)
$P_{\text{c}}[k],P_{\text{d}}[k]$ $\geq 0$ $\forall k=0,\ldots,K-1,$ (12d)
will satisfy (1) for all $N$ elements under PSC and is, thus, realizable.
@@
Eqs. (1a) and (3) imply (2a) and (1b) implies (2b). From Proposition 1, (12b) ensures satisfaction of (1c). Similarly, with $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$, Proposition 2 guarantees that (1d) holds, and the PSC with (12d) and (3) ensure that (1e) holds for all $N$ elements. Thus, any feasible composite dispatch is realizable.∎
Note that for a fixed composite battery, if $N$ increases, the cutting plane in (12b) approaches the cutting plane of the relaxed model. Furthermore, for $M\gg 1$, $\varepsilon$ in (9) tends to zero. That is, with faster control and more granular elements, (12) approaches the relaxed battery model defined in [Mazen_CDC_paper]; however, with guaranteed realizability.
@@ IV Simulation Results
We apply the realizable composite battery model (RCB) in (12) to minimize power tracking error for a given reference signal based on the objective function $\sum_{k=0}^{K-1}\left((P_{\text{c}}[k]-P_{\text{d}}[k])-P_{\text{ref}}[k]\right)^{2}$. For comparison, relaxed and robust models from [RobustNawaf] are used under balanced operation. Also, a Mixed-Integer Linear Program (MILP) formulation with unequal power-sharing for $N$ elements in (1) is used for comparison. The chosen element battery parameters are obtained from a Tesla Powerwall and we set $N=1000$ and $M=1$. The results from using each model are shown in Fig. 3.
@@
The RCB model and relaxed battery model produced similar optimal dispatch. However, this optimal dispatch was not realizable under equal power sharing (relaxed) causing the battery to saturate, resulting in the sudden drop in power seen in Fig. 3. The MILP, robust, and RCB models produced realizable results, with the robust formulation resulting in worse tracking performance. The MILP and RCB formulations produced comparable results, however, with a significant computational advantage for the RCB model.
@@ V Conclusion
A new linear dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements to realize simultaneous charging and discharging. A priority-based disaggregation strategy is described and used to characterize sufficient conditions on the composite battery dispatch under which element constraints are guaranteed to be satisfied. Previous battery optimization implicitly assumed equal power sharing between all elements. By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models that yield realizable power trajectories. Practical implementation of such dispatch strategies could also justify the use of linear models for composite storage systems.